OpenStudy (anonymous):

A store has a total of 15 bicycles and tricycles. There are 37 wheels. How many tricycles does the store have? medal will be awarded

OpenStudy (anonymous):

@Android17

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (michele_laino):

If I call with \(b\) the number of bicycles and with \(t\) the number of tricycles, then I can model your problem with this algebraic system: \[\left\{ \begin{gathered} b + t = 15 \hfill \\ 2b + 3t = 37 \hfill \\ \end{gathered} \right.\] please solve that system with respect to \(b\) and \(t\)

OpenStudy (anonymous):

help me on one more plz